Grade 11Mechanics

a particle is moving with uniform acceleration travels 24 m and 64m in 1st successive intervals of 4second each its initial velocity is/

rubiks cuber
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
 
let initial velocity u
and acceleration is a so from equation of motion
4u+8a=24
4u+16a+8a=64
a=40/16
u=1m/s
 
 
 
Regards
Arun (askIITians forum expert)